b94h:

Marcus spent 10 hours doing his homework last week. This week he spent 11 hours doing homework. He says that he spent 110% more time doing homework this week. How did he solve that

b94h:

Having a hard time understanding this

jimthompson5910:

Subtract the time values: 11-10 = 1 Divide that difference over the original value: 1/10 = 0.10 = 10% If he said he spent 10% more time doing hw, then he would be correct. However, 110% is incorrect. The formula I used was C = [ (B-A)/A ] * 100% where A = old value B = new value C = percent change

jimthompson5910:

Notice how 10% of 10 is equal to 0.10*10 = 1 So this means he does 10% extra work (aka 1 hour) and the baseline reference is from the perspective of the initial time value (10 hours)

b94h:

Ok well but how does 0.10*10=1 when in the zero place there is no number

jimthompson5910:

I'm not sure what you mean by "zero place"